# DiffCut: Catalyzing Zero-Shot Semantic Segmentation with Diffusion Features and Recursive Normalized Cut

### v2

@@ 4.4 Model Analysis
In this section, we assess the impact of hyperparameters $\tau$ and $\alpha$ over the segmentation performance. We report in Fig. 5 the mIoU for various $\alpha$ values, with respect to partitioning threshold values $\tau$ ranging from 0.3 to 0.97 on Cityscapes validation set. As $\alpha$ increases, we observe a dual effect. First, since a greater $\alpha$ value shrinks the affinity matrix components towards 0, the partitioning cost corresponding to the NCut value decreases, explaining the shift of the optimal threshold between the different curves. Second, as $\alpha$ increases, the range of $\tau$ values for which the method yields competitive performance widens, contributing to the overall robustness of the method. For example, DiffCut outperforms our own competitive baseline AutoSC for any $\tau$ between 0.35 and 0.67 when $\alpha=10$, whereas it only surpasses it between 0.92 and 0.96 when $\alpha=1$.  Qualitatively, we observe in Fig. 6 that as $\tau$ increases, the method uncovers finer segments in images.
@@
Our chosen diffusion backbone uses a UNet-based architecture which consists of an encoder $\mathcal{E}$, bottleneck $\mathcal{B}$ and decoder $\mathcal{D}$. The hierarchical features of the encoder, with spatial resolution of $128\times 128$, $64\times 64$ and $32\times 32$ respectively, are injected into the decoder $\mathcal{D}$ via skip connections. Considering the final self-attention modules at resolution $64\times 64$ and $32\times 32$ in the encoder and decoder, we demonstrate in Tab. 3, that the encoder’s features extracted at the lowest spatial resolution retain the most semantic information and are sufficient to reach optimal performance. In addition, combining different hierarchical features does not lead to any improvements and adds up to the computational burden.
@@
In the optic to extend DiffCut to an open-vocabulary setting, we propose in Tab. 4, a straightforward extension to associate a semantic label to each segmentation mask. Our proposed extension reaches competitive performance, even outperforming most baselines dedicated to the task of open-vocabulary zero-shot semantic segmentation. To do so, we leverage a frozen convolutional CLIP backbone and perform a mask-pooling over its semantic features to borrow its pretrained open-vocabulary recognition ability. Once mask proposals are predicted, they are classified with category text embeddings in a contrastive manner, where the class embeddings for each mask and category text embeddings are projected into a common embedding space.
@@ 5 Discussion
In this work, we address the challenging task of unsupervised and zero-shot image segmentation by introducing DiffCut, which significantly narrows the performance gap with fully supervised models. Specifically, DiffCut leverages diffusion features of a UNet encoder in a recursive graph partitioning algorithm to produce segmentation maps and reaches state-of-the-art performance on popular benchmarks. Yet, fully supervised end-to-end segmentation approaches still have an edge in terms of both performance and efficiency. Additional developments to further bridge the gap with supervised models could pave the way to use diffusion UNet encoders as foundation vision encoders for future downstream tasks.
@@ Acknowledgement
We thank Louis Serrano, Adel Nabli, and Louis Fournier for their insightful discussions and helpful suggestions on the paper.
@@ Appendix A Recursive Normalized Cut on Diffusion Features
We summarize in Algorithm 1 the recursive clustering process used in DiffCut.
@@
  • Raise to the power $\alpha$ each element of matrix $\bm{A}$ to obtain the affinity matrix $\bm{W}$ (see Eq. 3).
@@ Appendix C Additional Comparison with MaskCut
DiffCut, unlike MaskCut [32], is capable of providing dense segmentation maps and dynamically adapting the number of detected segments based on the visual content of an image. In contrast, MaskCut can only detect a fixed number of segments, making it less suitable for image segmentation. This limitation arises from MaskCut’s use of an iterative graph partitioning approach, where graph nodes associated with detected objects are masked. As a result, each segment is treated as a single object and cannot be refined after detection, which severely restricts its ability to identify a large number of objects. To highlight the superiority of our recursive partitioning over MaskCut’s iterative process, we present below a comparison between DiffCut and MaskCut with $k$ the number of objects to be detected varying in $\{3,5,20\}$.
@@
DiffCut significantly outperforms MaskCut, regardless of the chosen value of $k$. DiffCut’s improvement showcases our two key contributions: the effectiveness of our visual features for semantic segmentation and the ability of the recursive NCut algorithm to dynamically adjust the number of segments based on each image’s visual content.
@@ Appendix D DiffCut with Alternative Diffusion Backbones
To further display the relevance of diffusion features, we show that DiffCut achieves competitive performance even when using smaller diffusion backbones than SSD-1B. Specifically, we test two alternatives: SD1.4 and SSD-Vega [35] (another distilled version of SDXL). The UNet encoder in SD1.4 has 260M parameters, comprising approximately 30% of the overall UNet, while the UNet encoder in SSD-vega has 240M parameters, making up around 32% of the UNet.
@@
Model VOC Context COCO-Object COCO-Stuff-27 Cityscapes ADE20K SD1.4 57.5 52.8 30.0 45.2 24.5 36.7 SSD-Vega 62.2 56.4 34.9 49.5 30.1 45.7 SSD-1B 65.2 56.5 34.1 49.1 30.6 44.3 DiffSeg 49.8 48.8 23.2 44.2 16.8 37.7
The results obtained using these two backbones are consistent with those achieved with SSD-1B. While the SD1.4 UNet encoder shows a slight performance drop compared to SSD-1B, DiffCut still significantly outperforms DiffSeg. Notably, with the SSD-Vega UNet encoder, DiffCut delivers performance comparable to SSD-1B, despite having only half the number of parameters.
@@ Appendix E Mask Upsampling
Normalized Cut algorithm does not scale well with the graph size due to the generalized eigenvalue problem to solve, which hinder its use on the native image resolution (e.g., $1024\times 1024$). Thus, the clustering is applied in the latent space and yields segmentation maps at the latent resolution. To obtain pixel-level segmentation at the original image resolution, we need to upscale the low-resolution maps. In Tab. 8, we compare the nearest-neighbor upsampling approach versus our concept assignment upsampling and show that our proposed method obtain better results than the naive upsampling of the segmentation masks.
@@ Appendix F Visual Encoders KMeans Comparison
To evaluate the potential of vision encoders for zero-shot segmentation, we compare their clustering performance with a simple KMeans algorithm. For selected vision encoders, features are extracted from the last layer and clustered with KMeans whose hyperparameter $\bm{K}$ is either determined by the ground-truth ($\bm{K}^{*}$) for each image, or fixed across the dataset. We compute the mIoU with respect to the ground truth masks using the Hungarian matching algorithm [75]. Tab. 9 shows that the diffusion encoder (SSD-1B) significantly outperforms all other vision encoders on Pascal VOC (20 classes and no background), COCO-Stuff-27 and Cityscapes. This further confirms that the diffusion features are a good candidate for localizing and segmenting objects. Interestingly, unsupervised models such as DINOv2 are better than CLIP models, suggesting that text-aligned features does not contain accurate localization features.
@@ Appendix G Hyperparameter $\tau$ tuning
We show in Sec. 4.3 that the performance of the method is highly robust with respect to the value of the threshold $\tau$. However, as the granularity of annotations varies across target datasets, the value of this threshold, fixed in our experiments, can not be in the optimal regime on each benchmark. Therefore, we relax the condition on the absence of prior knowledge about the target dataset and report in Tab. 10 results of DiffCut where $\tau$ is loosely tuned using a small subset of annotated images (200) from the target training split. Specifically, we estimate an adequate value for $\tau$ with a grid-search in the set $\{0.35,0.55,0.75\}$ for COCO-Object, COCO-Stuff and Cityscapes.
@@
Appendix H Image Noising
Before passing the image to the diffusion UNet, a predefined amount of gaussian noise, controlled by a parameter called the timestep, is added to it. At timestep $t=0$ the input image corresponds to the original image without added noise while $t=1000$ corresponds to an image transformed into pure gaussian noise. In Fig. 7, we show the segmentation performance on the validation split of Pascal VOC for timesteps values ranging from 0 to 500. We can observe that a small amount of noise, around 50, gives the best mIoU score, indicating that the best semantic features are obtained with a slightly noisy input image. We note that despite a significant drop in the mIoU score for $t=500$, DiffCut still reaches state-of-the-art segmentation performance on Pascal VOC benchmark, demonstrating a high robustness of the method with respect to the noising ratio.
@@ Appendix H Image Noising
Appendix I Hyperparameter Sensitivity
We provide in Fig. 8 an additional evaluation of the combined effect of hyperparameters $\alpha$ and $\tau$ on the segmentation performance of DiffCut for Pascal VOC dataset. Similar to Fig. 5, we observe a shift of the optimal threshold between the different curves associated to distinct $\alpha$ values. We also observe the increased robustness of DiffCut with a wide range of $\tau$ values for which the method yields mIoU scores above 60.0 for $\alpha=10$. In comparison, for $\alpha=1$, the mIoU only surpasses 60.0 between 0.91 and 0.94.
@@ Appendix I Hyperparameter Sensitivity
Appendix J Visualization of the effect of $\tau$
@@
Appendix K Semantic Coherence in Vision Encoders
@@
Appendix L Additional Visualization
@@
Appendix M Datasets Licenses

### v3

@@ 4.4 Model Analysis
In this section, we assess the impact of hyperparameters $\tau$ and $\alpha$ over the segmentation performance. We report in Fig.˜5 the mIoU for various $\alpha$ values, with respect to partitioning threshold values $\tau$ ranging from 0.3 to 0.97 on Cityscapes validation set. As $\alpha$ increases, we observe a dual effect. First, since a greater $\alpha$ value shrinks the affinity matrix components towards 0, the partitioning cost corresponding to the NCut value decreases, explaining the shift of the optimal threshold between the different curves. Second, as $\alpha$ increases, the range of $\tau$ values for which the method yields competitive performance widens, contributing to the overall robustness of the method. For example, DiffCut outperforms our own competitive baseline AutoSC for any $\tau$ between 0.35 and 0.67 when $\alpha=10$, whereas it only surpasses it between 0.92 and 0.96 when $\alpha=1$. Qualitatively, we observe in Fig.˜6 that as $\tau$ increases, the method uncovers finer segments in images.
@@
Our chosen diffusion backbone uses a UNet-based architecture which consists of an encoder $\mathcal{E}$, bottleneck $\mathcal{B}$ and a decoder $\mathcal{D}$. The hierarchical features of the encoder, with spatial resolution of $128\times 128$, $64\times 64$ and $32\times 32$ respectively, are injected into the decoder $\mathcal{D}$ via skip connections. Considering the final self-attention modules at resolution $64\times 64$ and $32\times 32$ in the encoder and decoder, we demonstrate in Tab.˜3, that the encoder’s features extracted at the lowest spatial resolution retain the most semantic information and are sufficient to reach optimal performance. In addition, combining different hierarchical features does not lead to any improvements and adds to the computational burden.
@@
To extend DiffCut to an open-vocabulary setting, we propose in Tab.˜4, a straightforward approach that assigns a semantic label to each segmentation mask. After mask proposals are generated, an image is processed by a frozen convolutional CLIP visual encoder, which produces visual representations aligned with text in a shared embedding space via a projection layer. The embedding of each predicted segment is obtained by mask-pooling CLIP visual features, allowing a classification against category text embeddings through contrastive matching. Specifically, let $\mathbf{e}$ represent the embedding of a segment, and let $\{t_{i}\}_{i=1}^{N}$ denote the text embeddings of category names generated by the pretrained text encoder, the predicted class for this segment is determined as follows: $c=\arg\max_{i\in\{1,\cdots,N\}}\;softmax([cos(\mathbf{e},t_{1}),cos(\mathbf{e},t_{2}),\cdots,cos(\mathbf{e},t_{N})])$. This proposed extension reaches competitive performance, even outperforming several baselines dedicated to the task of open-vocabulary zero-shot semantic segmentation.
@@ 5 Discussion
In this work, we tackle the challenging task of unsupervised zero-shot semantic segmentation by introducing DiffCut, a method that significantly narrows the performance gap with fully supervised models. DiffCut leverages the diffusion features of a UNet encoder within a recursive graph partitioning algorithm to generate sharp segmentation maps, achieving state-of-the-art results on popular benchmarks. By reusing pre-trained models in a zero-shot manner, our approach can not only reduce computational resources, energy consumption, and human labor but also align with sustainable AI practices. However, because the diffusion backbone is not specifically trained for specialized domains, such as biomedical imaging, the method may struggle with out-of-distribution images. Fine-tuning the diffusion model on domain-specific data could mitigate this challenge. While fully supervised, end-to-end segmentation methods currently offer better efficiency and accuracy, further advancements could close this gap, positioning diffusion-based UNet encoders as foundation models for future vision tasks.
@@ Acknowledgement
We thank Louis Serrano, Adel Nabli, and Louis Fournier for their insightful discussions and helpful suggestions on the article. This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011014763 made by GENCI. We acknowledge the financial support provided by ANRT for its funding through the CIFRE grant 2022/0817 and PEPR Sharp (ANR-23-PEIA-0008, ANR, FRANCE 2030).
@@ Appendix A Recursive Normalized Cut on Diffusion Features
We summarize in Algorithm˜1 the recursive clustering process used in DiffCut.
@@
  • Raise to the power $\alpha$ each element of matrix $\bm{A}$ to obtain the affinity matrix $\bm{W}$ (see Eq.˜3).
@@ Appendix C Additional Comparison with MaskCut
DiffCut, is capable of providing dense segmentation maps and dynamically adapting the number of detected segments based on the visual content of an image. In contrast, MaskCut [32] can only detect a fixed number of segments, making it less suitable for image segmentation. This limitation arises from the use of an iterative graph partitioning approach, where graph nodes associated with detected objects are masked. Each segment is treated as a single object and cannot be refined after detection, which severely restricts its ability to identify a large number of objects. To highlight the superiority of our recursive partitioning over MaskCut’s iterative process, we present below a comparison between the two methods with $k$ the number of objects to be detected varying in $\{3,5,20\}$.
@@
DiffCut significantly outperforms MaskCut, regardless of the chosen value of $k$. DiffCut’s improvement shows our two key contributions: the effectiveness of our visual features for semantic segmentation and the ability of the recursive NCut algorithm to dynamically adjust the number of segments based on the visual content of each image.
@@ Appendix D DiffCut with Alternative Diffusion Backbones
To further display the relevance of diffusion features, we show that DiffCut achieves competitive performance even when using smaller diffusion backbones than SSD-1B. Specifically, we test two alternative models: SD1.4 and SSD-Vega [35] (another distilled version of SDXL). The UNet encoder in SD1.4 has 260M parameters, comprising approximately 30% of the overall UNet, while the UNet encoder in SSD-vega has 240M parameters, making up around 32% of the UNet.
@@
Model VOC Context COCO-Object COCO-Stuff-27 Cityscapes ADE20K DiffCut w/ SD1.4 57.5 52.8 30.0 45.2 24.5 36.7 DiffCut w/ SSD-Vega 62.2 56.4 34.9 49.5 30.1 45.7 DiffCut w/ SSD-1B 65.2 56.5 34.1 49.1 30.6 44.3 DiffSeg 49.8 48.8 23.2 44.2 16.8 37.7
The results obtained using SD1.4 and SSD-Vega are consistent with those achieved with SSD-1B. While the SD1.4 UNet encoder shows a slight performance drop compared to SSD-1B, DiffCut still significantly outperforms DiffSeg. Notably, with the SSD-Vega UNet encoder, DiffCut delivers performance comparable to SSD-1B, despite having only half the number of parameters.
@@ Appendix E Mask Upsampling
Normalized Cut algorithm does not scale well with the graph size due to the generalized eigenvalue problem to solve, which hinder its use on the native image resolution (e.g., $1024\times 1024$). Thus, the clustering is applied in the latent space and yields segmentation maps at the latent resolution. To obtain pixel-level segmentation at the original image resolution, we need to upscale the low-resolution maps. In Tab.˜8, we compare the nearest-neighbor upsampling approach versus our concept assignment upsampling and show that our proposed method obtain better results than the naive upsampling of the segmentation masks.
@@ Appendix F Visual Encoders KMeans Comparison
To evaluate the potential of vision encoders for zero-shot segmentation, we compare their clustering performance with a simple KMeans algorithm. For selected vision encoders, features are extracted from the last layer. The hyperparameter $\bm{K}$ is either determined by the ground-truth ($\bm{K}^{*}$) for each image, or fixed across the dataset. We compute the mIoU with respect to the ground truth masks using the Hungarian matching algorithm [75]. Tab.˜9 shows that the diffusion UNet encoder (SSD-1B) significantly outperforms other vision encoders on Pascal VOC (20 classes and no background), COCO-Stuff-27 and Cityscapes. This confirms that diffusion features are well-suited for localizing and segmenting objects. Interestingly, unsupervised models such as DINOv2 are better than CLIP models, suggesting that text-aligned features does not contain accurate localization features.
@@ Appendix G Hyperparameter $\tau$ tuning
We show in Sec.˜4.3 that the performance of the method is highly robust with respect to the value of the threshold $\tau$. However, as the granularity of annotations varies across target datasets, the value of this threshold, fixed in our experiments, can not be in the optimal regime on each benchmark. Therefore, we relax the condition on the absence of prior knowledge about the target dataset and report in Tab.˜10 results of DiffCut where $\tau$ is loosely tuned using a small subset of annotated images (200) from the target training split. Specifically, we estimate an adequate value for $\tau$ with a grid-search in the set $\{0.35,0.55,0.75\}$ for COCO-Object, COCO-Stuff and Cityscapes.
@@
Appendix H Hungarian Matching
Given a set of predicted masks, our goal is to find the best matching ground-truth labels. For each predicted mask, we compute the Intersection over Union (IoU) with every ground-truth mask and select the one with the highest IoU as its optimal ground-truth match. This results in a pairing where each predicted mask is associated with at most one ground truth mask. In datasets that include a background class, this label implicitly encompasses a variety of concepts related to "things" or "stuff," which vary depending on the dataset. Since our method generates a segment for every detected object in an image, a one-to-one matching between a single predicted cluster and the entire background does not accurately represent the model’s true categorization capabilities. Therefore, in such cases, we use a many-to-one matching approach by associating the clusters that primarily overlap with the background to its ID.
Appendix I Image Noising
Before passing the image to the diffusion UNet, a predefined amount of gaussian noise, controlled by a parameter called the timestep, is added to it. At timestep $t=0$ the input image corresponds to the original image without added noise while $t=1000$ corresponds to an image transformed into pure gaussian noise. In Fig.˜7, we show the segmentation performance on the validation split of Pascal VOC for timesteps values ranging from 0 to 500. We can observe that a small amount of noise, around 50, gives the best mIoU score, indicating that the best semantic features are obtained with a slightly noisy input image. We note that despite a significant drop in the mIoU score for $t=500$, DiffCut still reaches state-of-the-art segmentation performance on Pascal VOC benchmark, demonstrating a high robustness of the method with respect to the noising ratio.
@@ Appendix I Image Noising
Appendix J Hyperparameter Sensitivity
Fig.˜8 presents an additional evaluation of how the hyperparameters $\alpha$ and $\tau$ influence the segmentation performance on Pascal VOC. Similar to the observations in Fig.˜5, we notice a shift in the optimal threshold across the various curves corresponding to different $\alpha$ values. Besides, DiffCut shows increased robustness across a wide range of $\tau$ values, achieving mIoU scores exceeding 60.0 when $\alpha=10$. In contrast, for $\alpha=1$, the mIoU only exceeds 60.0 for $\tau$ between 0.91 and 0.94.
@@
Appendix K Runtime Comparison
Tab.˜11 presents a runtime comparison between DiffCut, MaskCut, and DiffSeg, which are the main baselines for graph-based image clustering and diffusion-based zero-shot segmentation, respectively.
Table 11: Runtime Comparison.
MaskCut $(k=5)$ DiffCut DiffSeg - SD1.4 DiffSeg - SSD-1B images / sec 0.84 1.11 2.75 1.25
Appendix L Visualization of the effect of $\tau$
@@
Appendix M Semantic Coherence in Vision Encoders
@@
Appendix N Additional Visualization
@@
Appendix O Datasets Licenses